OpenStudy (vshiroky):

The healthcare provider ordered amiodarone (Nexterone) to infuse at 0.8 mg/min. The pharmacy delivered amiodarone 150 mg in 100 mL D5W. At what rate will the nurse set the infusion? ___ mL/hr

OpenStudy (tkhunny):

Multiply by 1 \(\dfrac{0.8\;mg}{1\;min}\cdot\dfrac{3600\;min}{1\;hr}\cdot\dfrac{100\;mL\;D5W}{150\;mg}\) This is a wonderfully useful technique. Very clean. Very straight-forward. "Dimensional Analysis". Learn it. Use it.